OpenStudy (anonymous):

Solve using the addition principle. -23 = m + 18

OpenStudy (anonymous):

You can add 23 to both sides of the equation, so that you end up with: \[0 = m + 41\] Then, you can subtract the 41 from the right side to give you:\[m = -41\] I'm not sure if that's what your teacher means by the "addition principle". The simpler way would be to just subtract 18 from both sides to cancel out the 18 on the right side, and be done in one step.

jhonyy9 (jhonyy9):

-23=m+18 add to both sides (-18) -23+(-18)=m+18+(-18) -23-18=m+18-18 -41=m